Circular No. 8400

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 R1 (McNAUGHT) J. Young (Table Mountain, 0.6-m reflector + CCD) reports that images

taken in poor seeing (and at low altitude) on Sept. 6.13–6.16 UT show a coma diameter of ∼ 5′′. Additional CCD images by R. H. McNaught with the Siding-Spring 1.0-m reflector on Sept. 6.5 show an ill-defined center of brightness that is elongated in the direction of tail (and of the comet’s motion). Additional astrometry, the following preliminary orbital elements, and an ephemeris appear on MPEC 2004-R24.

T = 2004 Aug. 30.185 TT  ω = 0.^{o}573
e = 0.67898  Ω = 296.131 }2000.0
q = 0.98775 AU  i = 4.875
a = 3.07694 AU n^{o} = 0.182610 P = 5.40 years

SUPERNOVA 2002hh IN NGC 6946 M. Barlow and J. Fabbri, University College, London; and M. Meixner

and B. Sugerman, Space Telescope Science Institute, report the detection of the type-II supernova 2002hh in archival Spitzer IRAC images of NGC 6946 obtained by the SINGS Legacy program on 2004 June 10.76 UT. A source was detected in all four IRAC bands at α = 20^{h}34^{m}44^{s}.23, δ = +60^{o}07′19′′.5 (equinox 2000.0; ± 0′′.3 in each coordinate), in close agreement with the optical (IAUC 8005), radio (IAUC 8018), and x-ray positions (IAUC 8024). The measured flux densities are 0.93 ± 0.13 mJy at 3.6 μm, 1.24 ± 0.11 mJy at 4.5 μm, 4.22 ± 0.44 mJy at 5.8 μm, and 17.40 ± 1.85 mJy at 8.0 μm. This rise is fitted by a 300-K blackbody with integrated flux 1.35 × 10^{−11} erg s^{−1} cm^{−2}, which for a distance to NGC 6946 of 6 Mpc corresponds to a luminosity of 5.8 × 10^{40} erg s^{−1}. The equivalent blackbody radius of 10^{17} cm is too large to have been reached by dust-condensate ejecta in the 590 days since the supernova discovery date. Meikle et al. (IAUC 8024) derived a host-galaxy extinction of A(V) ∼ 5.0 to the supernova. A significant fraction of this extinction may arise from a pre-existing dust shell that was ejected during an earlier phase of evolution of the massive progenitor star and heated by the supernova event. Three convolved companion sources were also detected, with the following position angles, separations, and fluxes (mJy) at 3.6, 4.5, 5.8, and 8.0 μm: Star 2 (= 2MASS 20344320+6007234), 297^{o}, 8′′.6, 81.4 ± 9.0, 62.6 ± 3.5, 55.0 ± 1.9, 36.9 ± 1.9; Star 3, 286^{o}, 11′′.4, 3.5 ± 1.6, 0.6 ± 0.1, 0.8 ± 0.3, 2.1 ± 0.4; Star 4, 319^{o}, 3′′.2, 1.0 ± 0.4, 0.6 ± 0.1, 3.6 ± 0.6, 12.8 ± 1.8.

2004 September 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green